OpenStudy (anonymous):

What are the solutions of the quadratic equation? x^2+11x=-24

OpenStudy (jdoe0001):

hint: \(x^2+11x=-24\implies \begin{array}{cccllll} x^2&+11&+24=0\\ &\uparrow &\uparrow \\ &8+3&8\cdot 3 \end{array}\)

OpenStudy (anonymous):

ok thank you